Find the (a) mean, (b) median, (c) mode, and (d) midrange for the given sample data. An experiment was conducted to determine whether a deficiency of carbon dioxide in the soil affects the phenotype of peas. Listed below are the phenotype codes where 1 3-wrinkled-yellow, and 4 wrinkled-green. Do the results make sense? 4 smooth-yellow, 2-smooth-green, 3 2 2 (a) The mean phenotype code is (Round to the nearest tenth as needed.) (b) The median phenotype code is (Type an integer or a decimal.) (c) Select the correct choice below and fill in any answer boxes within your choice. O A. The mode phenotype code is (Use a comma to separate answers as needed.) O B. There is no mode (d) The midrange of the phenotype codes is (Type an integer or a decimal.) Do the measures of center make sense? O A. Only the mean, median, and mode make sense since the data is numerical. O B. All the measures of center make sense since the data is numerical. O C. Only the mode makes sense since the data is nominal.
Answer #1

Solution:

Mean =(4+1+3+3+3+2+2+3+1+3+1+2+2+3)/14 = 33/14 = 2.36

For median let's sort the data from smallest to largest

1,1,1,2,2,2,2,3,3,3,3,3,3,4

So median is (14+1)/2 = 7.5th value

So median is (2+3)/2=2.5

Mode for this data is 3

Midrange for this data is

(Largest-Smallest)/2 = (4-1)/2 =1.5

Also its answer is C. I.e. only mode makes sense since the data is nominal.